# Efficient Diversity-based Experience Replay for Deep Reinforcement Learning

### v2

@@ B.4 Ablation Experiments
Trajectory length $b$ is a key parameter in reinforcement learning tasks as it determines how much information the agent can gather from the environment. To investigate the impact of this parameter on learning performance, we designed a series of experiments by varying the trajectory length $b$ and observing its effect on the model’s performance.In the experiments, the trajectory length $b$ was set to 2, 5, 10, and 20, and EDER’s performance was evaluated in different test environments. Longer trajectories typically provide more comprehensive information about the environment’s dynamics, aiding strategy learning, but also increase computational complexity. The results show that a moderate trajectory length (e.g., $b=10$) performs best in most cases, achieving a good balance between learning efficiency and computational burden. Figure  8 presents the learning curves of EDER with different trajectory lengths, further validating this conclusion.
@@ C.1 Environment Setup
Mujoco Environment.The Mujoco environment is a widely-used physics engine for robot control and continuous action space tasks. It simulates realistic physical interactions, making it ideal for testing strategies in complex manipulation tasks and fine motor control. In our experiments, we focused on two classic task environments: FetchEnv and HandEnv, representing different types of robotic manipulation tasks, and conducted multiple experiments to evaluate the performance of the EDER method in these environments.The picture can be seen in Figuare  11

### v4

@@ 6 Conclusion
Acknowledgments
This work was supported in part by the Key Research and Development Project of Hubei Province (2022BCA057), the Science and Technology Development Fund Macau SAR (0003/2023/RIC, 0052/2023/RIA1, 0031/2022/A, 001/2024/SKL for SKL-IOTSC), the Shenzhen-Hong Kong-Macau Science and Technology Program Category C (SGDX20230821095159012), the National Natural Science Foundation of China (62402325), and the Research Foundation of Shenzhen Polytechnic University (6022310014K, 6022312054K). This work was performed in part on the supercomputing system at the Supercomputing Center of Wuhan University and in part at SICC which is supported by SKL-IOTSC, University of Macau.
@@ B.4 Ablation Experiments
Trajectory length $b$ is a key parameter in reinforcement learning tasks as it determines how much information the agent can gather from the environment. To investigate the impact of this parameter on learning performance, we designed a series of experiments by varying the trajectory length $b$ and observing its effect on the model’s performance.In the experiments, the trajectory length $b$ was set to 2, 5, 10, and 20, and EDER’s performance was evaluated in different test environments. Longer trajectories typically provide more comprehensive information about the environment’s dynamics, aiding strategy learning, but also increase computational complexity. The results show that a moderate trajectory length (e.g., $b=10$) performs best in most cases, achieving a good balance between learning efficiency and computational burden. Figure  7 presents the learning curves of EDER with different trajectory lengths, further validating this conclusion.
@@ C.1 Environment Setup
Mujoco Environment.The Mujoco environment is a widely-used physics engine for robot control and continuous action space tasks. It simulates realistic physical interactions, making it ideal for testing strategies in complex manipulation tasks and fine motor control. In our experiments, we focused on two classic task environments: FetchEnv and HandEnv, representing different types of robotic manipulation tasks, and conducted multiple experiments to evaluate the performance of the EDER method in these environments.The picture can be seen in Figuare  10